# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

### v1

@@ 5.4.1. Engine-Agnostic Processing
In Data-Juicer 1.0, the data processing pipeline is implemented using two distinctive execution modes: a standalone mode tailored for single-node operations and a distributed mode designed for scalability over multiple nodes. These modes exploit different dataset classes, each with a unique set of functionalities and interfaces. The default standalone execution mode employs the Hugging Face Dataset class (Lhoest et al., 2021), which provides a rich array of encapsulated functions, such as dataset.map() and dataset.filter(). It’s also equipped with configurable batch processing capabilities essential for various computational needs. In contrast, the distributed execution mode leverages the Ray Dataset class (Moritz et al., 2018), which scales effectively across multiple nodes.
@@
Despite both dataset classes using a storage format based on Apache Arrow (Arrow, 2023), they exhibit significant differences in the behaviors and interfaces exposed. For instance, the Ray Dataset delineates individual and batched sample processing using separate methods: map and map_batches. It also allows to specify GPU counts for optimized scheduling. Meanwhile, the Hugging Face Dataset excels in supporting a broad spectrum of data modalities, such as image and audio. Moreover, the Hugging Face Dataset is typically applied in read-heavy data processing scenarios, such as in-memory tokenization and tensor reshaping, which are crucial for training deep learning models. For processing tasks in the context of foundation models, especially those involving synthesis operations, write-heavy procedures warrant attention. Here, the Ray Dataset provides flexible data exportation techniques advantageous to such tasks.
@@
As highlighted in Sec. 5.1, we introduce a top-level DJ-Dataset class in Data-Juicer 2.0, along with common functions to bridge the variety of interfaces and implementation specifics across diverse computational engines. Besides the support of Hugging Face Dataset and Ray Dataset, our DJ-Dataset is also seamlessly extendable to support distributed computing within Alibaba Cloud ecosystems, thanks to the compatibility between ours and MaxFrame-DataFrame classes with intermediate in-memory formats like Pandas (pandas development team, 2020) and NumPy (Harris et al., 2020) or external-memory formats like Parquet. Thus, we develop unified wrappers that extract the core execution functions of different DJ-Operators into User-Defined Functions for the MaxCompute engine. Compared to Spark, MaxCompute is compatible in both syntax and runtime on Alibaba Cloud nodes and can be considered a commercially optimized version. An internal empirical comparison shows that MaxCompute SQL achieves 50% better performance than native Spark SQL.
@@
The Deduplicators within Data-Juicer 2.0. Fuzzy deduplication is complex, involving a mixture of operations such as map, filter, group by, aggregate, and join (Broder et al., 2000). The support and performance of these operations vary across different engines, especially in large-scale scenarios. To demonstrate the engine-agnostic feature of Data-Juicer 2.0, we use our minhash_deduplicator as an example, which supports the aforementioned three different engines. Users only need to specify algorithm-specific parameters such as jaccard_threshold and num_permutations. These parameters remain consistent regardless of the engine used, while engine-specific details and optimizations are completely transparent to the user. For instance, we utilize Ray Actors to implement our Ray-based deduplicator, starting with a load-balanced, distributed union-find algorithm (Kim et al., 2024) and introducing a hash-based aggregation process to enhance memory utilization and efficiency. This method avoids fragmented unions caused by Ray’s native groupby operation, which is computationally expensive in typical LSH implementations with traditional big-data engines (Mou, 2023). As a result, we achieve a 3.3x speedup over our vanilla Ray version. More performance evaluations are presented in Sec. 6.1.4.
5.4.2. Fault Tolerance
@@
Figure 5. Illustration of fault tolerance for “bad” data.
@@
5.4.3. Streaming I/O and Subset Splitting
@@
Secondly, we develop a user-friendly script that automatically pre-splits the original dataset based on two observations: (1) The size limit of the underlying Apache-Arrow block, and (2) The inherent automatic block-splitting strategy in Ray. When there are tens of thousands of nodes but with only a few dataset files, Ray would split the dataset files according to the available resources and distribute the blocks of the dataset to all nodes, which brings a huge network communication cost and decreases the CPU utilization of each node. Thus, we split the original dataset into smaller 128MB files in advance automatically according to the dataset size and the number of distributed nodes, trying to adapt the features of Arrow and Ray for better performance (an empirical study can be found in Sec. 6.1.4). This approach reduces location and reprocessing costs associated with fault tolerance and helps mitigate network exchange overheads, especially beneficial in contexts involving large-scale multimodal data, as well as in scenarios that require handling global objects of Ray Actor in distributed modes.
5.5. Internal Adaptation for DJ-Operators
This section explores the internal adaptations developed for DJ-Operators, which are crucial for optimizing resource allocation and user experience without requiring users to understand hardware specifics or implementation details. We implement several automatic adaptation features for resource management, aiming at balancing resource constraints and operational efficiency within Data-Juicer 2.0. These strategies include automatic OP reordering at the recipe level (Sec. 5.5.1), automatic parallelism at the OP level (Sec. 5.5.2), and automatic data insight mining to assess the impact of each OP on data samples, considering both upstream and downstream OPs (Sec. 5.5.3). These features are encapsulated in a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime, with default sample size as min(1000, remaining_data_size).
5.5.1. Workloads-Aware OP Reordering
@@
(1) $v_{fused}=\frac{N}{T_{total}}=\frac{N}{\sum_{i=1}^{n}\frac{N}{v_{i}}}=\frac{1}{\sum_{i=1}^{n}\frac{1}{v_{i}}},$
where $N$ is the total number of samples to be processed by the fused OP and $T_{total}$ is the total time cost of it. Then, faster OPs are prioritized, while slower ones are deferred to later stages. This probe-based approach identifies the globally optimal reordering solution for each OP group, outperforming the suboptimal strategies of the previous version, as empirically validated in Sec. 6.2.1.
5.5.2. Automatic Operator-Wise Parallelism
@@
  • Model-based OPs that integrate substantial models require considerable computational time, potentially spanning hundreds of hours on CPUs for large datasets. To mitigate this, Data-Juicer 2.0 expedites these OPs by automatically leveraging CUDA and GPUs when available. Given the diverse memory requirements of large models, we utilize the Adapter component to conduct quick VRAM benchmarking prior to full-size dataset processing. This information is systematically assigned to the mem_required parameter of the respective OPs. To further alleviate GPU memory demand, we also integrate quantization libraries such as vLLM (Kwon et al., 2023) to enable efficient inference of leveraged models.
@@
  • For general-purpose and I/O-intensive OPs, Data-Juicer 2.0 enables batched processing, using a robust default batch_size parameter guided by performance saturation trends shown in Sec. 6.2.3. Batched processing reduces I/O overhead, boosts efficiency per processor, and enhances overall speed. Additionally, we introduce hierarchical parallelism for OPs involving multimodal data I/O, such as image_aspect_ratio_filter. These OPs utilize multi-process and GPU parallelism, along with multiple threads, to handle data batches more efficiently, taking the concurrent opportunities between I/O and computation latencies.
5.5.3. Automatic Operator-Wise Insight Mining
The cumulative impact of a sequence of OPs on a dataset can be complex, and the combined effectiveness is not always greater than that of a single OP. This is demonstrated by experiments on cutting-edge foundation models for text and video (Chen et al., 2024c). Traditionally, systems like Data-Juicer 1.0 (Chen et al., 2024a) and Falcon (Penedo et al., 2023) primarily generated coarse-grained visualizations such as Sankey diagrams to show percentage changes in data volume, mainly through Filters.
Figure 6. Illustration of OP-wise insight mining.
To formulate better data recipes, we introduce fine-grained OP-level insight mining and visualization, extending beyond Sankey diagrams. As shown in Fig. 6, our system provides comprehensive tracking of data statistics (e.g., stats.perplexity) and category tags (e.g., meta.text_category_tags). After the processing completion of each OP, we use the Adapter’s probing function along with Data-Juicer’s embedded Analyzer. Built-in category tagging OPs across modalities and analytic Filters help obtain histograms of interested semantic categories and relevant statistics.
Data-Juicer 2.0 uses these analytics to automatically assess changes in metrics between successive OPs, generating detailed visualization reports that highlight significant differences along lineages. This reporting aids in analyzing each OP’s significance, understanding correlations, and optimizing data recipes. For instance, if an image-text matching Filter (using a model like BLIP-2 (Li et al., 2023a)) reduces text length statistics significantly, it may indicate noisy, too-long captions that could affect textual description accuracy for text-image pairs.
6. System Experiments
In this section, we present practical patterns and anti-patterns of Data-Juicer 2.0 based on extensive performance evaluations.
6.1. Cloud-scale Data Processing Efficiency
6.1.1. Experiment Settings.
@@ 6.1. Cloud-scale Data Processing Efficiency
We start with a base dataset comprising 560k image-text pairs (about totally 5.6 million textual tokens). We expand this dataset by factors of {2, 4, 20, 100, 500, 2500, 12500, 125000}, resulting in nine datasets scaling to 70B data samples. These datasets are categorized into three scales: small (1x, 2x, 4x), medium (4x, 20x, 100x), and large (100x, 500x, 2500x, 12500x, 125000x). For different scales, we prepare both multimodal and text-only data processing recipes, each containing five OPs. We run the recipes on these datasets using various computing engines (Standalone, Ray, MaxCompute) with different Alibaba Cloud resources, including ECS instances, PAI-DLC, and MaxCompute, spanning 1 to 100 nodes and 64 to 12,800 CPU cores. To ensure a fair comparison, we use the same worker node configuration, such as CPU frequency, across different computing engines.
(a) Small-scale multimodal datasets.
(b) Small-scale text-only datasets.
(c) Medium-scale multimodal datasets.
(d) Medium-scale text-only datasets.
(e) Large-scale text-only datasets.
(f) Resource utilization comparison.
Figure 7. Data processing time comparison with different computing engines, dataset sizes, and workloads.
6.1.2. The Case of Small Scales
As shown in Fig. 7a, when processing small-scale multimodal datasets, the standalone mode with Hugging Face Dataset is efficient and comparable to the Ray mode with a single node. Additional Ray nodes provide further but limited acceleration (speedup ratios between 138% and 226% with 4 nodes). For text-only datasets (Fig. 7b), the standalone mode remains efficient. However, for the Ray mode, 4x node increments yield smaller speedups (148%) or even increased processing times due to dominated I/O and communication costs. Thus, with Data-Juicer, processing datasets with hundreds of thousands of samples on a single machine is both efficient and cost-effective for most users.
6.1.3. The Case of Medium Scales.
When datasets are scaled to 56M samples, processing times increase significantly with the standalone mode, thus, its performance is omitted in Figures 7c and 7d. Here, the Ray mode outperforms in all instances, demonstrating considerable speedups with increasing node counts, making it the recommended choice for medium-scale scenarios. Moreover, compared to native ECS nodes (green lines), the Ray mode on PAI-DLC (red lines) is faster, saving 24.8% of processing time due to the Alibaba Cloud’s optimization dedicated to cluster networking.
6.1.4. The Case of Large Scales.
For datasets at the 70B-sample scale, professional cloud-scale distributed data processing products are advantageous. Users benefit from the vast cloud computing resources without dealing with the intricacies of setup and management. From our experiments, we recommend Ray-DLC for multimodal recipes and MaxCompute for text-only recipes at this scale. (1) For multimodal recipes, using 3200 Ray-DLC cores process datasets in 1780.86s and 7083.5s for 500x and 2500x dataset sizes, respectively, indicating good scalability. On the other hand, the MaxCompute engine requires 1.5 times more processing time using the same resources, due to the challenges of loading large-size multimodal data. (2) For text-only recipes (Fig. 7e), although the Ray mode benefits from additional cores even at the scale of ten thousand, MaxCompute is the fastest, requiring about 1/4 of the time and using 1/2 fewer cores. This is attributed to MaxCompute’s co-optimization of distributed computing and storage, a feature not as advanced in Ray’s implementation.
Impact of I/O and Data Splitting. I/O plays a crucial role in data processing workloads. For distributed processing, large-scale datasets need to be distributed to nodes, stored on shared disks (e.g., NAS), or on cloud storage services (e.g., Alibaba Cloud OSS and CPFS). All these introduce unavoidable I/O and network communication costs. In our experiments shown in Fig. 7, we use Alibaba Cloud standard CPFS, finding that NAS and OSS can increase time costs by 20-30% compared to CPFS. Furthermore, with another AI-oriented CPFS product that has a 3x larger network bandwidth and optimized RDMA, processing the 12500x dataset with 3200 cores in Ray mode takes 1625s, 2.7 times faster than the standard CPFS version (4396s). We then scale up the dataset to 125000x, resulting times as 14617s and 7611s with 3200 and 6400 cores respectively, further demonstrating the system’s efficiency and scalability.
Additionally, the adaptive subset splitting capability for Ray mode, introduced in Sec. 5.4.3, offers 2$\sim$3x speedups in our experiments. For example, with 100 Ray-DLC nodes and 12800 cores, our subset splitting method reduces the processing time of 12500x dataset from over 5000s to about 2000s. Analysis of Ray’s resource utilization shows that its automatic block-splitting strategy incurs high network communication costs and low CPU utilization when many nodes handle a few dataset files. Our strategy pre-splits datasets according to size and node count, optimizing alignment with Ray’s features and enhancing efficiency. As seen in Fig. 7f, this strategy reduces peak network I/O from about 160MB/s to 60MB/s and achieves more consistent CPU utilization across all nodes, thus accelerating dataset processing.
Table 1. Deduplication time (minutes) with Ray nodes across varying dataset sizes and CPU counts (160 cores per node).
{tblr}
4*160 11.13 min 50.83 min 285.43 min
8*160 7.47 min 30.08 min 168.10 min
Large-scale Deduplication. We tested the MinHash-based RayDeduplicator with datasets from Redpajama (Computer, 2023) and Common Crawl chunks (Crawl, 2023), sized at 200GB, 1TB, and 5TB, using CPU counts ranging from 640 to 1280 cores. As shown in Table 1, RayDeduplicator efficiently scales with increased data size and computing resources, indicating its capability to handle large-scale deduplication tasks effectively. When the data size increases by 5x, the processing time increases by 4.02$\sim$5.62x. When the number of CPU cores doubles, the processing time decreases to 58.9$\sim$67.1of the original time. Notably, RayDeduplicator can process 5TB of data in 2.8 hours using 8 Ray nodes (8*160 CPU cores). In contrast, NVIDIA’s NeMo Curator, which leverages cuDF (cuDF Team, 2024) and Dask (Team, 2024a), takes 1.8 hours to process 1.1TB of Redpajama data using 64 A100 GPUs (64*6192 CUDA cores), as shown in (Team, 2024b).
6.2. Ablation Study on Runtime Adaptations
6.2.1. Automatic Workloads-aware OP Reordering.
@@ 6.2. Ablation Study on Runtime Adaptations
We evaluate the performance improvements due to our probe-based OP reordering on two recipes with different numbers of OPs. The simple recipe contains 5 OPs, with 2 fusible OPs, while the complex recipe includes 13 OPs, with 5 fusible OPs. We run these recipes on the base dataset used before (580k image-text pairs), comparing processing times for all OPs, fusible OPs, and other OPs.
@@
Figure 8. Processing time for simple (left) and complex (right) recipes. Processing times for different types of OPs are scaled according to the time of baseline without OP fusion.
6.2.2. Automatic GPU Resource Allocation
Efficient model inference on GPUs is critical for the processing speeds of model-based OPs. We select four image-related OPs, including an image-only OP and three image-text cross-modality OPs, integrating various models that require different GPU memories. Using a test dataset of 10k image-text pairs, we process these OPs on a machine with 64 processors and compare the times with and without GPU support in Data-Juicer 2.0. Table 2 shows these comparative results.
Table 2. Multimodal OPs processing time comparison between CPU cores and GPUs on 10k samples. “np” denotes the number of processors.

### v2

@@ F.1 Engine-agnostic Processing
In Data-Juicer 1.0, the data processing pipeline is implemented using two distinctive execution modes: a standalone mode tailored for single-node operations and a distributed mode designed for scalability over multiple nodes. These modes exploit different dataset classes, each with a unique set of functionalities and interfaces. The default standalone execution mode employs the Hugging Face Dataset class [36], which provides a rich array of encapsulated functions, such as dataset.map() and dataset.filter(). It’s also equipped with configurable batch processing capabilities essential for various computational needs. In contrast, the distributed execution mode leverages the Ray Dataset class [45], which scales effectively across multiple nodes.
@@
Despite both dataset classes using a storage format based on Apache Arrow [3], they exhibit significant differences in the behaviors and interfaces exposed. For instance, the Ray Dataset delineates individual and batched sample processing using separate methods: map and map_batches. It also allows to specify GPU counts for optimized scheduling. Meanwhile, the Hugging Face Dataset excels in supporting a broad spectrum of data modalities, such as image and audio. Moreover, the Hugging Face Dataset is typically applied in read-heavy data processing scenarios, such as in-memory tokenization and tensor reshaping, which are crucial for training deep learning models. For processing tasks in the context of foundation models, especially those involving synthesis operations, write-heavy procedures warrant attention. Here, the Ray Dataset provides flexible data exportation techniques advantageous to such tasks.
@@
As highlighted in Sec. 4.1, we introduce a top-level Data-Juicer-Dataset class in Data-Juicer 2.0, along with common functions to bridge the variety of interfaces and implementation specifics across diverse computational engines. Besides the support of Hugging Face Dataset and Ray Dataset, our Data-Juicer-Dataset is also seamlessly extendable to support distributed computing within Alibaba Cloud ecosystems, thanks to the compatibility between ours and MaxFrame-DataFrame classes with intermediate in-memory formats like Pandas [48] and NumPy [30] or external-memory formats like Parquet. Thus, we develop unified wrappers that extract the core execution functions of different Data-Juicer-Operators into User-Defined Functions for the MaxCompute engine. Compared to Spark, MaxCompute is compatible in both syntax and runtime on Alibaba Cloud nodes and can be considered a commercially optimized version. An internal empirical comparison shows that MaxCompute SQL achieves 50% better performance than native Spark SQL.
@@
The Deduplicators within Data-Juicer 2.0. Fuzzy deduplication is complex, involving a mixture of operations such as map, filter, group by, aggregate, and join [9]. The support and performance of these operations vary across different engines, especially in large-scale scenarios. To demonstrate the engine-agnostic feature of Data-Juicer 2.0, we use our minhash_deduplicator as an example, which supports the aforementioned three different engines. Users only need to specify algorithm-specific parameters such as jaccard_threshold and num_permutations. These parameters remain consistent regardless of the engine used, while engine-specific details and optimizations are completely transparent to the user. For instance, we utilize Ray Actors to implement our Ray-based deduplicator, starting with a load-balanced, distributed union-find algorithm [34] and introducing a hash-based aggregation process to enhance memory utilization and efficiency. This method avoids fragmented unions caused by Ray’s native groupby operation, which is computationally expensive in typical LSH implementations with traditional big-data engines [46]. As a result, we achieve a 3.3x speedup over our vanilla Ray version.
F.2 Fault Tolerance
@@
Figure 7: Illustration of fault tolerance for “bad” data.
@@
F.3 Streaming I/O and Subset Splitting
@@
Secondly, we develop a user-friendly script that automatically pre-splits the original dataset based on two observations: (1) The size limit of the underlying Apache-Arrow block, and (2) The inherent automatic block-splitting strategy in Ray. When there are tens of thousands of nodes but with only a few dataset files, Ray would split the dataset files according to the available resources and distribute the blocks of the dataset to all nodes, which brings a huge network communication cost and decreases the CPU utilization of each node. Thus, we split the original dataset into smaller 128MB files in advance automatically according to the dataset size and the number of distributed nodes, trying to adapt the features of Arrow and Ray for better performance. This approach reduces location and reprocessing costs associated with fault tolerance and helps mitigate network exchange overheads, especially beneficial in contexts involving large-scale multimodal data, as well as in scenarios that require handling global objects of Ray Actor in distributed modes.
Appendix G Implementation of Adaptation for Data-Juicer-Operator
This section explores the internal adaptations developed for Data-Juicer-Operator, which are crucial for optimizing resource allocation and user experience without requiring users to understand hardware specifics or implementation details. We implement several automatic adaptation features for resource management, aiming at balancing resource constraints and operational efficiency within Data-Juicer 2.0. These strategies include automatic OP reordering at the recipe level (Appendix G.1), automatic parallelism at the OP level (Appendix G.2), and automatic data insight mining to assess the impact of each OP on data samples, considering both upstream and downstream OPs (Appendix G.3). These features are encapsulated in a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime, with default sample size as min(1000, remaining_data_size).
G.1 Workloads-aware OP Reordering
@@
$v_{fused}=\frac{N}{T_{total}}=\frac{N}{\sum_{i=1}^{n}\frac{N}{v_{i}}}=\frac{1}{\sum_{i=1}^{n}\frac{1}{v_{i}}},$ (1)
where $N$ is the total number of samples to be processed by the fused OP and $T_{total}$ is the total time cost of it. Then, faster OPs are prioritized, while slower ones are deferred to later stages. This probe-based approach identifies the globally optimal reordering solution for each OP group, outperforming the suboptimal strategies of the previous version, as empirically validated in Appendix H.2.1.
G.2 Automatic Operator-wise Parallelism
@@
  • Model-based OPs that integrate substantial models require considerable computational time, potentially spanning hundreds of hours on CPUs for large datasets. To mitigate this, Data-Juicer 2.0 expedites these OPs by automatically leveraging CUDA and GPUs when available. Given the diverse memory requirements of large models, we utilize the Adapter component to conduct quick VRAM benchmarking prior to full-size dataset processing. This information is systematically assigned to the mem_required parameter of the respective OPs. During the execution of these OPs on datasets, Data-Juicer 2.0 continuously assesses the available GPU memory of the execution environment to dynamically and adaptively determine the optimal parallelism strategy. To further alleviate GPU memory demand, we also integrate quantization libraries such as vLLM [35] to enable efficient inference of leveraged models.
@@
  • For general-purpose and I/O-intensive OPs, Data-Juicer 2.0 enables batched processing, using a robust default batch_size parameter guided by performance saturation trends shown in Appendix H.2.3. Batched processing reduces I/O overhead, boosts efficiency per processor, and enhances overall speed. Additionally, we introduce hierarchical parallelism for OPs involving multimodal data I/O, such as image_aspect_ratio_filter. These OPs utilize multi-process and GPU parallelism, along with multiple threads, to handle data batches more efficiently, taking the concurrent opportunities between I/O and computation latencies.
G.3 Insight Mining Example
Here we provide an example of OP-wise insight mining in Fig. 8.
Figure 8: Illustration of OP-wise insight mining.
Appendix H Additional Experiment Details and Results
H.1 Implementation Details of Scalability Experiments
@@
We start with a base dataset comprising 560k image-text pairs (about totally 5.6 million textual tokens) from the pertaining dataset of LLaVA [40]. We expand this dataset by factors of {2, 4, 20, 100, 500, 2500, 12500, 125000}, resulting in nine datasets scaling to 70B data samples. These datasets are categorized into three scales: small (1x, 2x, 4x), medium (4x, 20x, 100x), and large (100x, 500x, 2500x, 12500x, 125000x). For different scales, we prepare both multimodal and text-only data processing recipes, each containing 5 OPs. We run the recipes on these datasets using various computing engines (Standalone, Ray, MaxCompute) with different Alibaba Cloud resources, including ECS instances, PAI-DLC, and MaxCompute, spanning 1 to 100 nodes and 64 to 12,800 CPU cores. To ensure a fair comparison, we use the same worker node configuration, such as CPU frequency, across different computing engines.
H.2 Ablation Study on Runtime Adaptations
H.2.1 Automatic workloads-aware OP reordering
@@
We evaluate the performance improvements due to our probe-based OP reordering on two recipes with different numbers of OPs. The simple recipe contains 5 OPs, with 2 fusible OPs, while the complex recipe includes 13 OPs, with 5 fusible OPs. We run these recipes on the base dataset used before (560k image-text pairs), comparing processing times for all OPs, fusible OPs, and other OPs.
@@
Figure 9: Processing time for simple and complex recipes.
H.2.2 Automatic GPU resource allocation
Efficient model inference on GPUs is critical for the processing speeds of model-based OPs. We select four image-related OPs, including an image-only OP and three image-text cross-modality OPs, integrating various models that require different GPU memories. Using a test dataset of 10k image-text pairs, we process these OPs on a machine with 64 processors and 8 Nvidia A100 GPUs, and compare the times with and without GPU support in Data-Juicer 2.0. Table 2 shows these comparative results.
Table 2: Multimodal OPs processing time comparison between CPU cores and GPUs on 10k samples. “np” denotes the number of processors.